OpenStudy (anonymous):

A number divided by 4 is 3 more than –11. What is the number? What is the correct solution to this problem? –2 –32 –56 1 points

OpenStudy (texaschic101):

n/4 = -11 + 3 n/4 = - 8 --- multiply both sides by 4 n = -8 * 4 n = - 32 check... -32/4 = -11 + 3 -8 = -8 (correct) so your number (n) = -32

OpenStudy (anonymous):

thx helped me alot